use the properties of logarithm to write each expression as a single logarithm with a coefficient of 1. ln(3)-2ln(4+8),ln(3)-2(ln(4)+ln(8),ln(3)-2ln(4)+ln8
Answer by lwsshak3(11628)

use the properties of logarithm to write each expression as a single logarithm with a coefficient of 1. ln(3)-2ln(4+8),ln(3)-2(ln(4)+ln(8),ln(3)-2ln(4)+ln8
**
doing 3 expressions separately
ln(3)-2ln(4+8)=ln[3/12^2]
ln(3)-2(ln(4)+ln(8)=ln[24/4^2]
ln(3)-2ln(4)+ln8=(same as previous expression)